# Top Score on the Wrong Exam: On Benchmarking in Machine Learning for Vulnerability Detection

### v1

@@ IV Empirical Study Design
Figure 6: Function Size: Lines of code of the 100 functions per dataset that we sampled for our empirical study.
@@
    RQ.2 Can we achieve a high classification performance even when the root cause of the vulnerability is hidden? How can we explain the excellent performance of ML4VD on these widely-used benchmarks despite this severe flaw in the problem statement? Do the popular datasets contain properties that ML4VD techniques can exploit to achieve high accuracy without actually detecting vulnerabilities?
IV-A Methodology
@@
Selection Criteria. In order to select datasets that represent the state-of-the-art of ML4VD benchmarks, we chose the top-most widely-used datasets from the software engineering literature (via our literature survey) for which the authors provide links to patch commits and/or CVE websites (required for manual labeling). Additionally, we reviewed unpublished literature for emerging datasets and included all publicly available datasets with more than 50 citations in 2024.
Selected datasets. Based on our selection criteria, we chose the BigVul [38], Devign [39], and DiverseVul [23] datasets. The selection of BigVul and Devign resulted from our literature survey, as these datasets were the most popular among the 22 ML4VD papers we analyzed. However, we excluded ReVeal [40] from our study since the authors did not publish patch commit IDs or CVEs, which are necessary for determining vulnerability and context dependence. Additionally, we included the DiverseVul dataset [23], a recently published dataset (2023) that has seen significant use in many yet unpublished papers based on our review of unpublished literature. Figure 4 illustrates the citations of the selected datasets measured by Google Scholar on July 31, 2024. All three datasets are becoming increasingly popular, with each having more than 50 citations in 2024 alone.
Sampling. From each of the three datasets, we randomly selected 100 samples labeled as vulnerable, using a sample size inspired by related studies on data quality, such as the one by Croft et al., which used a sample size of 70 [33]. We have published the reproducible script and the sampled functions as part of our artifact, which is available at https://github.com/niklasrisse/TopScoreWrongExam. The resulting 300 functions come from patch commits published between 2010 and 2022, covering 80 unique open-source projects (BigVul: 25, Devign: 2, DiverseVul: 60). Figure 6 illustrates the size of these functions across the three datasets. While we did not formally categorize the vulnerability types, we frequently observed issues such as out-of-bounds writes/reads, improper restriction of operations within memory buffer bounds, improper input validation, and use-after-free vulnerabilities.
Labeling Process. In a time-intensive process (more than 120 hours), one of the authors of our paper (a software engineering researcher) reviewed each of the 300 functions. The process involved opening the corresponding patch commit on GitHub to understand the vulnerability. If available, we also reviewed the CVE (Common Vulnerabilities and Exposures) report for additional context. For full transparency, we release all labels and explanations we generated as part of our artifact available at https://github.com/niklasrisse/TopScoreWrongExam.
Figure 5 visualizes the complete process that we employed for each individual function. In all cases where the patch commit was available, and a decision could be made within 15 minutes, we assigned one of the following labels:
  • $\star$
  • $\star$
Additionally, we added a short explanation in natural language to justify the assigned label, which we used to analyze the results more in-depth. To ensure the accuracy of this first labeling step, 200 out of the 300 functions were independently cross-labeled by another software engineering researcher. The two raters achieved a Cohen Kappa value of 0.7 [41] (85% of agreement), which implies substantial agreement according to the guidelines provided by Landis and Koch [42]. All cases of disagreement were resolved through discussion and resulted in a single label/explanation for each function.
For functions labeled as vulnerable (indicating they actually were the source of the vulnerability), we conducted a second round of labeling to determine context dependency:
  • $\star$
  • $\star$
Again, we provided a short explanation in natural language to justify the context label.
Reproducibility. To ensure the reproducibility of our empirical study and to provide transparency in our research, we have made all related scripts and data publicly available. This includes all of the intermediate and final results of our empirical study. All resources can be accessed at https://github.com/niklasrisse/TopScoreWrongExam.
V Results
RQ.1-a Noisy Labels
@@ V Results
BigVul 9 (16%) 29 (51%) 19 (33%)
Devign 31 (68%) 14 (30%) 1 (2%)
DiverseVul 1 (3%) 20 (67%) 9 (30%)
Results. Figure 7 shows the results of our empirical study. For RQ.1-a), we are interested in the proportion of functions labeled as vulnerable that actually contain security vulnerabilities, which are displayed in the left part of each subfigure (vulnerable vs. secure). Out of the 100 functions per dataset that were originally labeled as vulnerable, we found only 38%-64% to actually contain security vulnerabilities. 30%-57% do not contain security vulnerabilities and were therefore labeled to be secure. For 2%-4%, we were not able to make a decision after 15 minutes, and for 1%-4%, we could find neither the patch commit nor the CVE report (not available).
Noisy Labels (RQ1-a). Out of the 100 functions per dataset that were originally labeled as vulnerable, only 38%-64% (Devign: 50%, BigVul: 38%, DiverseVul: 64%) actually contain security vulnerabilities.
Based on our evidence, we can confirm the findings of Croft et al. [33]. The differences in label accuracy, especially for the Devign dataset (80% accurate labels found by them vs. 50% found by us), might be explained by two reasons: First, Croft et al. considered label accuracy for all functions, and secure functions are more likely correctly labeled. In RQ.1-a, we only establish label accuracy for functions labeled as vulnerable. Second, Croft et al. establish label accuracy more conservative than us, where a function can be considered correctly labeled as vulnerable if it calls a function known to be vulnerable. In this case, we decided that only the actual vulnerable function should be labeled as vulnerable because the calling function would not be vulnerable if the actual vulnerable function would be fixed.
@@
BigVul 16 (42%) 19 (50%) 1 (3%) 2 (5%) 0 (0%)
Devign 23 (47%) 21 (43%) 0 (0%) 5 (10%) 0 (0%)
DiverseVul 25 (39%) 34 (53%) 2 (3%) 1 (2%) 2 (3%)
Patch Commit Identification. The first reason for label inaccuracy are errors during the process of identifying patch commits. From the original papers [39, 23, 38], we know that all three datasets start their data collection process by identifying vulnerability-patching commits in popular open-source software repositories. However, the Devign dataset identifies these commits only by filtering the commit messages for security-related keywords. We observe that 68% of falsely labeled functions in our sample of the Devign dataset originate from this automatic identification process. The triangulation via other data sources seems to address this source of inaccuracy for the other datasets.
@@ 34 GFP_KERNEL);
(a) Dependence on external functions: The heap-based buffer overflow (CVE-2019-14816) in lines 20-21 of this function from the BigVul dataset depends on the external function cfg80211_find_vendor_ie. Without knowing this external function, we do not know what values vs_ie can have, and consequently, we do not know whether the buffer overflow can ever be triggered.
1PP_Flash_Menu* ReadMenu(int depth,
2 const IPC::Message* m,
3 PickleIterator* iter) {
4 if (depth > kMaxMenuDepth)
5 return NULL;
6 ++depth;
7
8 PP_Flash_Menu* menu = new PP_Flash_Menu;
9 menu->items = NULL;
10
11 if (!m->ReadUInt32(iter, &menu->count)) {
12 FreeMenu(menu);
13 return NULL;
14 }
15
16 if (menu->count == 0)
17 return menu;
18
19 menu->items = new PP_Flash_MenuItem[menu->count];
20PP_Flash_Menu* ReadMenu(int depth,
21PP_Flash_Menu* ReadMenu(int depth,
22 const IPC::Message* m,
23 PickleIterator* iter) {
24 if (depth > kMaxMenuDepth)
25 return NULL;
26 ++depth;
27
28 PP_Flash_Menu* menu = new PP_Flash_Menu;
29 menu->items = NULL;
30
31 if (!m->ReadUInt32(iter, &menu->count)) {
32 FreeMenu(menu);
33 return NULL;
34 }
35
36 if (menu->count == 0)
37 return menu;
38
39 menu->items = new PP_Flash_MenuItem[menu->count];
41 for (uint32_t i = 0; i < menu->count; ++i) {
42 if (!ReadMenuItem(depth, m, iter, menu->items + i)) {
43 FreeMenu(menu);
44 return NULL;
45 }
46 }
47 return menu;
48}
(b) Dependence on type declaration: The integer overflow (CVE-2013-0892) in line 20 of this function from the BigVul dataset depends on PP_Flash_MenuItem, menu->items and menu->count. Only if sizeof(PP_Flash_MenuItem) * menu->count exceeds the range of menu->items, the operation overflows.
@@ 43 FreeMenu(menu);
16 if (!(newrealm = find_realm_data(sprinc->realm.data,
17 (krb5_ui_4) sprinc->realm.length)))
@@
(c) Dependence on function argument: The null pointer dereference (CVE-2013-6800) in line 8 of this function from the BigVul dataset depends on the function argument sprinc. Only if sprinc=NULL it crashes with a null pointer dereference in line 8.
@@
(d) Dependence on globals: Line 7 of this function is the cause of multiple buffer overflows in the JasPer repository (CVE-2014-8158). Whether line 7 is actually compiled depends on whether the macro HAVE_VLA is defined.
Figure 8: Examples for context-dependent security vulnerabilities.
Structural Changes. The second reason for label inaccuracy is structural changes. All three datasets included in our study assume that all functions changed by a vulnerability-patching commit were vulnerable before the patch was applied. However, according to our results, only a subset of the functions changed by a patch commit is actually vulnerable, while other functions could be changed to address structural changes that are a consequence of fixing the actual vulnerable function. For instance, fixing a buffer overflow may require adding a buffer size parameter to the function call wherever the function is called. This can lead to false labels if all functions changed by the patch commit are considered to be vulnerable before the patch was applied. In fact, 30%-67% of falsely labeled functions can be attributed to structural changes.
Unrelated Changes. The third reason for label inaccuracy are other unrelated changes to functions in vulnerability patch commits. These include stylistic changes, e.g. removing whitespace or adding comments. According to our study, 2%-33% of falsely labeled functions can be attributed to such unrelated changes.
RQ.1-b Context-Dependent Vulnerability
Results. Each of the actual vulnerable functions that resulted from the first step of our empirical study (RQ.1 (a)) was assigned one of two labels: Context-independent or context-dependent. Figure 7 show the results of this second labeling round. To our surprise, all 152 vulnerable functions in our study required additional context to be accurately identified (context-dependent). Not a single function could be detected without considering any additional context beyond the function itself (context-independent)
Context Dependence (RQ1-b). All 152 vulnerable functions in our empirical study required additional context to be identified. Without this additional context, it was impossible to determine whether these functions actually contain security vulnerabilities.

### v2

@@ IV Empirical Study Design
    RQ.2 Can we achieve a high classification performance even when the root cause of the vulnerability is hidden? How can we explain the excellent performance of ML4VD on these widely-used benchmarks despite this severe flaw in the problem statement? Do the popular datasets contain properties that ML4VD techniques can exploit to achieve high scores without actually detecting vulnerabilities?
IV-A Methodology RQ.1
Selection Criteria. In order to select datasets that represent the state-of-the-art of ML4VD benchmarks, we chose the top-most widely-used datasets from the ML4VD literature (via our literature survey) for which the authors provide links to patch commits and/or CVE websites (required for manual labeling). Additionally, we reviewed unpublished literature for emerging datasets (pre-prints from arXiv777arXiv: https://arxiv.org).
Selected datasets. Based on our selection criteria, we chose the BigVul [29], Devign [30], and DiverseVul [1] datasets. The selection of BigVul and Devign resulted from our literature survey, as these datasets were the most popular among the 81 ML4VD papers we analyzed. However, we excluded ReVeal [31] from our study since the authors did not publish patch commit IDs or CVEs, which are necessary for determining vulnerability and context dependence. Additionally, we included the DiverseVul dataset [1], a recently published dataset (2023) that has seen significant use in many yet unpublished papers based on our review of unpublished literature. Figure 4 illustrates the citations of the selected datasets measured by Google Scholar on January 16, 2025. All three datasets are becoming increasingly popular, with each having more than 100 citations in 2024 alone.
Sampling. From each of the three datasets, we randomly selected 100 samples labeled as vulnerable, using a sample size inspired by related studies on data quality, such as the one by Croft et al., which used a sample size of 70 [20]. We have published the reproducible script and the sampled functions as part of our GitHub repository, which is available at https://github.com/niklasrisse/TopScoreWrongExam. The resulting 300 functions come from patch commits published between 2010 and 2022, covering 80 unique open-source projects (BigVul: 25, Devign: 2, DiverseVul: 60). While we did not formally categorize the vulnerability types, we frequently observed issues such as out-of-bounds writes/reads, improper restriction of operations within memory buffer bounds, improper input validation, and use-after-free vulnerabilities.
Labeling Process. In a time-intensive process (more than 150 hours), two Software Security researchers (co-authors of this paper) reviewed each of the 300 functions independently. The process involved opening and understanding the corresponding patch commit on GitHub and, if available, reviewing the CVE (Common Vulnerabilities and Exposures) report for additional context.
Figure 5 visualizes the complete process that the two researchers employed for each individual function. In all cases where the patch commit was available, and a decision could be made within 15 minutes, they assigned one of the following labels:
  • $\star$
  • $\star$
Additionally, they added a short explanation in natural language to justify the assigned label, which we used to analyze the results more in-depth. For the first labelling step (secure or vulnerable), the labels of the two researchers agreed for 82% of all functions, resulting in a Cohen Kappa value of 0.64 [32], which implies substantial agreement according to the guidelines provided by Landis and Koch [33]. All cases of disagreement were resolved through discussion and resulted in a single label/explanation for each function. In the case that only one of the researchers was able to make a decision within 15 minutes, the researchers also entered a discussion for resolution, which resulted in a joint decision (vulnerable/secure) for all of these cases. Only if both researchers were unable to make a decision after 15 minutes, the final label was ”No decision”. In cases where the source of a vulnerability was ambiguous, e.g. if a vulnerability in a function could be fixed both by adding a check within the function (callee) or within a calling function (caller), the researchers only considered the callee to be vulnerable.
For functions labeled as vulnerable (indicating they actually were the source of the vulnerability), the two researchers conducted a second round of labeling to determine context dependency:
  • $\star$
  • $\star$
Again, the researchers provided a short explanation in natural language to justify the context label. For the second labelling step (context dependence), the labels of the two researchers agreed for 98% of all functions, resulting in a Cohen Kappa value of 0.96 [32], which implies almost perfect agreement according to the guidelines provided by Landis and Koch [33]. Again, all cases of disagreement were resolved through discussion and resulted in a single label/explanation for each function.
Labeling Expertise. Both Software Security researchers that carried out the labeling process have considerable experience in the Software Security research field (C/C++ security, in particular). The first researcher has three years of experience in Software Security research, published multiple papers on the topic of vulnerability detection at the top Security and Software Engineering conferences, and worked as a Software Engineer in industry for two years. The second researcher has five years of experience in Software Security, published multiple pre-prints on the topic (currently under submission at top conferences), and participated in discovering three CVEs, which are actually part of at least one of the three datasets we investigated for the empirical study.
Reproducibility. To ensure the reproducibility of our empirical study and to provide transparency in our research, we have made all related scripts and data publicly available. All resources can be accessed as part of our GitHub repository, which is available at https://github.com/niklasrisse/TopScoreWrongExam.
IV-B Methodology RQ.2
In order to test whether the three widely-used datasets (Devign, BigVul, and DiverseVul) contain spuriously correlated features that can be exploited to achieve a high performance without actually detecting security vulnerabilities, we trained a simple classifier (Gradient Boosting Classifier) to detect vulnerabilities based on word counts only, completely disregarding the structure of the code. After training, we evaluated the resulting models on the evaluation subsets of the respective datasets, again based on word counts only.
Metrics. We selected the f1-score as our primary performance metric, since it is the dominant metric in the literature we surveyed for Section III. Additionally, we report accuracy for the Devign dataset, because it is a relatively balanced dataset (47.9% vulnerable functions). Since both BigVul (4.5% vulnerable functions) and DiverseVul (5.4% vulnerable functions) are heavily imbalanced towards functions labeled as secure, using accuracy as a performance metric for model comparison could be misleading [34]. For example, in a dataset with 95.5% of functions labeled as secure (such as BigVul), a technique can achieve an accuracy of 95.5% by classifying all functions as secure. This is why we chose to omit accuracy for BigVul and DiverseVul.
Selected Techniques. Based on the literature we surveyed for Section III, the state-of-the-art techniques for the Devign dataset are SNOPY [26] with 67.86% f1-score and PDBERT [35] with 67.61% accuracy. For the BigVul dataset, the state-of-the-art technique is DeepDFA [36] with 96.46% f1-score. According to the benchmark provided by Chen et al. [1], CodeT5 Small [37] is the best performing technique for the DiverseVul dataset with 48.28% f1-score.
V Results
RQ.1-a Noisy Labels
@@ V Results
Results. Figure 6 shows the results of our empirical study. We were able to find the patch commit for all functions included in the empirical study, which is why the label ”Not Available” was omitted from Figure 6. For RQ.1-a), we are interested in the proportion of functions labeled as vulnerable that actually contain security vulnerabilities, which are displayed in the left part of each subfigure (vulnerable vs. secure). Out of the 100 functions per dataset that were originally labeled as vulnerable, we found only 39%-65% to actually contain security vulnerabilities. 35%-61% do not contain security vulnerabilities and were therefore labeled to be secure. For 0%-3%, we were not able to make a decision after 15 minutes.
Noisy Labels (RQ1-a). Out of the 100 functions per dataset that were originally labeled as vulnerable, only 39%-65% (Devign: 47%, BigVul: 39%, DiverseVul: 65%) actually contain security vulnerabilities.
Based on our evidence, we can confirm the findings of Croft et al. [20]. The differences in label accuracy, especially for the Devign dataset (80% accurate labels found by them vs. 47% found by us), might be explained by two reasons: First, Croft et al. considered label accuracy for all functions, and secure functions are more likely correctly labeled. In RQ.1-a, we only establish label accuracy for functions labeled as vulnerable. Second, Croft et al. establish label accuracy more conservative than us, where a function can be considered correctly labeled as vulnerable if it calls a function known to be vulnerable. In this case, we decided that only the actual vulnerable function should be labeled as vulnerable because the calling function would not be vulnerable if the actual vulnerable function would be fixed.
Patch Commit Identification. The first reason for label inaccuracy are errors during the process of identifying patch commits. From the original papers [30, 1, 29], we know that all three datasets start their data collection process by identifying vulnerability-patching commits in popular open-source software repositories. However, the Devign dataset identifies these commits only by filtering the commit messages for security-related keywords. We observe that 54% of falsely labeled functions in our sample of the Devign dataset originate from this automatic identification process. The triangulation via other data sources (e.g. CVE databases) makes this reason occur less frequently for the other datasets (21% for BigVul and 3% for DiverseVul).
Structural Changes. The second reason for label inaccuracy is structural changes. All three datasets included in our study assume that all functions changed by a vulnerability-patching commit were vulnerable before the patch was applied. However, according to our results, only a subset of the functions changed by a patch commit is actually vulnerable, while other functions could be changed to address structural changes that are a consequence of fixing the actual vulnerable function. For instance, fixing a buffer overflow may require adding a buffer size parameter to the function call wherever the function is called. This can lead to false labels if all functions changed by the patch commit are considered to be vulnerable before the patch was applied. In fact, 38%-66% of falsely labeled functions can be attributed to structural changes.
@@
BigVul 13 (21%) 29 (48%) 19 (31%)
Devign 27 (54%) 19 (38%) 4 (8%)
DiverseVul 1 (3%) 23 (66%) 11 (31%)
Unrelated Changes. The third reason for label inaccuracy are other unrelated changes to functions in vulnerability patch commits. These include stylistic changes, e.g. removing whitespace or adding comments. According to our study, 8%-31% of falsely labeled functions can be attributed to such unrelated changes.
RQ.1-b Context-Dependent Vulnerability
Results. Each of the actual vulnerable functions that resulted from the first step of our empirical study (RQ.1 (a)) was assigned one of two labels: Context-independent or context-dependent. Figure 6 show the results of this second labeling round. To our surprise, all 151 vulnerable functions in our study required additional context to be accurately identified (context-dependent). Not a single function could be detected without considering any additional context beyond the function itself (context-independent)
Context Dependence (RQ1-b). All 151 vulnerable functions in our empirical study required additional context to be identified. Without this additional context, it was impossible to determine whether these functions actually contain security vulnerabilities.
@@
BigVul 16 (41%) 19 (49%) 1 (2%) 3 (8%) 0 (0%)
Devign 22 (47%) 20 (43%) 0 (0%) 5 (10%) 0 (0%)
DiverseVul 26 (40%) 34 (52%) 2 (3%) 1 (2%) 2 (3%)
Dependence on External Functions. The first and most prevalent type of dependence we identified is dependence on external functions. Consider the example in Figure 7. The heap-based buffer overflow in lines 20-21 of this function from the BigVul dataset depends on the external function cfg80211_find_vendor_ie. Without knowing this external function, we do not know what values vs_ie can have, and consequently, we do not know whether the buffer overflow can ever be triggered. In our empirical study, 43%-52% of context-dependent vulnerabilities can be attributed to dependence on external functions.
Dependence on Function Arguments. The second type of dependence we identified is dependence on function arguments. Consider the example in Figure 8. The null-pointer-dereference vulnerability in line 8 of this function from the BigVul dataset depends on the function argument sprinc. However, without knowing the context in which this function can be called, we do not know whether sprinc can ever be NULL. For example, it could be properly validated before passing it to the function in all cases where this function is actually called. In that case, there would be no vulnerability. In our empirical study, 40%-47% of context-dependent vulnerabilities can be attributed to dependence on function arguments.
Dependence on Type Declarations. The third type of dependence we identified is dependence on type declarations. Consider the example in Figure 9. The integer overflow vulnerability in line 20 depends on the type declaration of PP_Flash_MenuItem. Only if sizeof(PP_Flash_MenuItem) * menu->count exceeds the range of menu->items, the operation overflows. In our empirical study, 0-3% of context-dependent vulnerabilities can be attributed to dependence on type declarations.
Dependence on Globals. The fourth type of dependence we identified is dependence on globals, such as macros and global variables. Consider the example in Figure 10. Line 7 of this function from the Devign dataset is the cause of multiple buffer overflows in the JasPer repository (CVE-2014-8158). If HAVE_VLA is not defined and QMFB_JOINBUFSIZE is smaller than bufsize but still used without successful dynamic allocation, any attempt to use joinbuf for storing data will result in writing beyond its allocated size (QMFB_JOINBUFSIZE). In our empirical study, 2-10% of context-dependent vulnerabilities can be attributed to dependence on globals.
@@ 34 GFP_KERNEL);
Figure 7: Dependence on external functions: The heap-based buffer overflow (CVE-2019-14816) in lines 20-21 of this function from the BigVul dataset depends on the external function cfg80211_find_vendor_ie. Without knowing this external function, we do not know what values vs_ie can have, and consequently, we do not know whether the buffer overflow can ever be triggered.
@@
16 if (!(newrealm = find_realm_data(sprinc->realm.data, (krb5_ui_4) sprinc->realm.length)))
@@
24}
Figure 8: Dependence on function argument: The null pointer dereference (CVE-2013-6800) in line 8 of this function from the BigVul dataset depends on the function argument sprinc. Only if sprinc=NULL it crashes with a null pointer dereference in line 8.
1PP_Flash_Menu* ReadMenu(int depth, const IPC::Message* m, PickleIterator* iter) {
2 if (depth > kMaxMenuDepth)
3 return NULL;
4 ++depth;
5
6 PP_Flash_Menu* menu = new PP_Flash_Menu;
7 menu->items = NULL;
8
9 if (!m->ReadUInt32(iter, &menu->count)) {
10 FreeMenu(menu);
11 return NULL;
12 }
13
14 if (menu->count == 0)
15 return menu;
16
17 menu->items = new PP_Flash_MenuItem[menu->count];
19 for (uint32_t i = 0; i < menu->count; ++i) {
20PP_Flash_Menu* ReadMenu(int depth, const IPC::Message* m, PickleIterator* iter) {
21PP_Flash_Menu* ReadMenu(int depth, const IPC::Message* m, PickleIterator* iter) {
22 if (depth > kMaxMenuDepth)
23 return NULL;
24 ++depth;
@@ 10 FreeMenu(menu);
26 PP_Flash_Menu* menu = new PP_Flash_Menu;
27 menu->items = NULL;
28
29 if (!m->ReadUInt32(iter, &menu->count)) {
30 FreeMenu(menu);
31 return NULL;
32 }
33
34 if (menu->count == 0)
35 return menu;
36
37 menu->items = new PP_Flash_MenuItem[menu->count];
39 for (uint32_t i = 0; i < menu->count; ++i) {
40 if (!ReadMenuItem(depth, m, iter, menu->items + i)) {
41 FreeMenu(menu);
42 return NULL;
43 }
44 }
45 return menu;
46}
Figure 9: Dependence on type declaration: The integer overflow (CVE-2013-0892) in line 20 of this function from the BigVul dataset depends on PP_Flash_MenuItem, menu->items and menu->count. Only if sizeof(PP_Flash_MenuItem) * menu->count exceeds the range of menu->items, the operation overflows.
@@ 41 FreeMenu(menu);
Figure 10: Dependence on globals: Line 7 of this function is the cause of multiple buffer overflows in the JasPer repository (CVE-2014-8158). Whether line 7 is actually compiled depends on whether the macro HAVE_VLA is defined.
RQ.1-c Context-dependent Security
Now we know that the vulnerability of all 151 vulnerable functions in our empirical study could not be decided without considering additional context, but what about secure functions? What proportion of functions labeled as secure—because they are not known to cause a vulnerability in the program—could potentially cause a vulnerability if an appropriate external context existed?
Methodology. From each of the three datasets (BigVul, Devign, and DiverseVul), we randomly selected 30 samples that were labeled as secure and checked if we could come up with a reasonable context for that function that would cause a vulnerability in the program. We note that we did not attempt to verify whether a function that was labeled as secure was actually not causing a vulnerability in the program. There can only be certainty about the presence of vulnerabilities (e.g., via a triggering input). In practice, there is no certainty about their absence. However, we publish the reproducible script and the sampled functions as part of our artifact for full transparency.
To find out what proportion of secure functions could be vulnerable if an appropriate external context would exist, we tried to construct an artificial vulnerable context for each of them. Consider the example in Figure 11. In this artificially crafted context, the function HTTP_Clone (marked in green) contains a null pointer dereference vulnerability. However, it is labeled as secure in the DiverseVul dataset because in its original context [^8] https://github.com/varnishcache/varnish-cache/commit/c5fd097e it is not known to contain a security vulnerability. Just looking at that function itself, without any external context, we cannot decide whether this function does not actually cause a vulnerability in the program. Similar to this example, we manually tried to construct a vulnerable context for all 90 functions in our sample. For full transparency, we include the generated vulnerable settings (as explanations in natural language) in our artifact.
RQ.1 Result Summary
Within our sample of functions, about half of those labeled as vulnerable did not actually contain the security vulnerabilities they were claimed to have. All of those that do are context-dependent vulnerabilities, i.e., these functions are vulnerable only because an appropriate context exists under which the function is vulnerable. Similarly, the majority of secure functions contain context-dependent vulnerabilities that would make that function vulnerable if an appropriate context existed. As there is evidently insufficient information in the base units—that are used for training, validation, and testing in these datasets—to decide their vulnerability without further context, we conclude that ML4VD cannot be soundly evaluated as a classic binary classification problem on the function-level.
Problem statement (RQ.1). ML4VD techniques cannot be soundly evaluated as a classic binary classification problem on intra-procedural base units, such as functions.